OpenStudy (anonymous):

two people shake hands once. Three people shake hands three times. i) How many times would 5 people shake hands ii) Try and find a general formula.

OpenStudy (dumbcow):

i) 10 times 1st guy shakes 4 other people's hands 2nd guy shakes remaining 3 peoples hands 3rd guy shakes remaining 2 hands 4th and 5th guys shake hands 4+3+2+1 = 10 ii) general formula if there are n people -->f(n) = (n-1) + (n-2) +...1 this forms an arithmetic sequence sum = n(n-1)/2

OpenStudy (dumbcow):

there is another way too are you learning permutations and combinations right now ?

OpenStudy (anonymous):

no just number patterns

OpenStudy (dumbcow):

ok then that answer is sufficient :)

OpenStudy (anonymous):

good grief you must be an genius

OpenStudy (dumbcow):

no not at all, just have taken a lot of classes...plus i like math

OpenStudy (anonymous):

lol thanks